# Leggett-Garg Inequality for Bosons in a Double-Well Potential

### v1

@@ Abstract
Figure 4: (a) Violation of the LGI. (b) Three correlation functions $C_{12}$ (the red dashed line), $C_{23}$ (the black solid line) and $C_{13}$ (the green dotted line) as functions of $\tau$. (c)/(d) Time evolution of $P(n_{\rm L})$ in the absence/presence of the measurement at $tJ/2\pi=0.2$ (represented by the gray dotted line). $n_{\rm L}=70$ at $tJ/2\pi=0.2$ in (d). (e) Colormap of $LG$ for all different initial occupation number in the left well $n_{\rm L}(0)$. The violation of the LGI is symmetric under the reflection with respect to the line $n_{\rm L}(0)=50$. The blue color in (e) shows the part where the LGI is violated. In (a)-(e), the parameters are $N=100$ and $\Lambda=0$.
Single particle regime.—Next, we study violation of the LGI in the absence of on-site interaction. In this case the Hamiltonian is simply given by $\hat{H}_{\rm BJJ}=-J\hat{S}_{x}$. For this model, the Schrödinger equation is analytically solvable and the resultant wave function at time $t$ evolved from the initial state $|N-l\rangle_{\rm L}|l\rangle_{\rm R}$ is given by
@@
The first/second line of Eq. (Leggett-Garg Inequality for Bosons in a Double-Well Potential) represents the coherent oscillation of the bosons initially located in the left/right well. These bosons flow in opposite directions. The former represents bosons initially tunneling from the left well to the right well, as shown in Fig. 4, while the latter represents those initially tunneling from the right well to the left well. As shown in Fig. 4, the peak of the occupation probability coherently oscillates between $n_{\rm L}=0$ and $n_{\rm L}=N$ without damping. Due to this coherent oscillation the LGI is violated, as shown in Fig. 4.
@@
Focusing on the first violation in Fig. 4, this behavior can be explained as follows: Until $C_{13}$ starts to decrease, the LGI holds, since $C_{12}=C_{13}=C_{23}=1$. The LGI begins to be first violated when $C_{13}$ starts decreasing at $\tau J/2\pi\simeq 0.125$. Since $C_{13}$ abruptly becomes $-1$ whilst $C_{12}=1$ and $C_{23}$ is still positive, the LGI becomes violated quite abruptly. The value of the peak of $LG$ becomes larger with the increase of $N$, due to more abrupt decrease of $C_{13}$ (see SM).
@@
The persistence of this violation is attributed to the bosons tunneling backwards in Fig. 4 induced by the measurement, which corresponds to the second line of Eq. (Leggett-Garg Inequality for Bosons in a Double-Well Potential). In the case of $C_{23}$ the bosons tunneling backwards impede the decrease of $C_{23}$, whilst such effect is absent in the case of $C_{13}$, where no measurement is performed at $t_{2}$ (see Fig. 4). The LGI again begins to hold approximately when $C_{12}$ becomes zero at $\tau J/2\pi\simeq 0.25$ (see Fig. 4).
@@
We also test the LGI for all different initial occupation numbers of the left well $n_{\rm L}(0)$ with $N=100$. In Fig. 4, it can be seen that the peak value of $LG$ increases with a larger initial population imbalance (see SM).
@@
Conclusion and discussion.—In this work, we study violations of the LGI for bosons in a double-well potential. We find that the LGI is violated due to Josephson oscillations in the weak coupling regime, while it remains unviolated in the strong coupling regime, attributed to the self-trapping phenomena. Remarkably, we observe that the LGI remains violated even in the large $N$ system. We also find that, in this system, on-site interaction induces a dephasing effect that destroys temporal correlations, resulting in the LGI no longer being violated afterwards. We have confirmed that the LGI is broken below the critical value of interaction strength for self-trapping. In contrast, it has been shown that the LGI is no longer violated when self-trapping takes place. Furthermore, we have revealed that the LGI is broken for nearly the same parameters for which the Josephson oscillations are observed in the experimental work of Ref. [19], thus indicating the presence of macroscopic quantum coherence in the semiclassical regime.
Our result suggests that either MRPS or NIM is violated, or possibly both, in the system of bosons in a double-well potential with a large number of particles. In our LG tests, we perform projection measurements, which necessarily affect the subsequent dynamics. Therefore, if measurements can be made without collapsing the state in any way [46], the result can be different from ours.
@@
Acknowledgments.—We thank G. Kimura, M. Kunimi and T. Nikuni for fruitful discussions. ST thanks F. Dalfovo, G. Ferrari, and A. Recati for useful discussions. RY is partially supported by a Grant-in-Aid of MEXT for Scientific Research (KAKENHI Grant Nos. 19K14616 and 20H01838). ST is supported by the Japan Society for the Promotion of Science Grant-in-Aid for Scientific Research (KAKENHI Grant No. 19K03691).

### v3

@@ Abstract
Figure 4: (a) Violation of the LGI. (b) Three correlation functions $C_{12}$ (the red dashed line), $C_{23}$ (the black solid line), and $C_{13}$ (the green dotted line) as functions of $\tau$. (c)/(d) Time evolution of $P(n_{\rm L})$ in the absence/presence of the measurement at $tJ/2\pi=0.2$ (represented by the gray dotted line). $n_{\rm L}=70$ at $tJ/2\pi=0.2$ in (d). (e) Color map of $K$ for all different initial occupation numbers in the left well $n_{\rm L}(0)$. The violation of the LGI is symmetric under the reflection with respect to the line $n_{\rm L}(0)=50$. The blue color in (e) shows the part where the LGI is violated. In (a)–(e), the parameters are $N=100$ and $\Lambda=0$.
@@
The first (second) line of Eq. (Violation of the Leggett-Garg inequality for dynamics of a Bose-Einstein condensate in a double-well potential) represents the coherent oscillation of the bosons initially located in the left (right) well. These bosons flow in opposite directions. The former represents bosons initially tunneling from the left well to the right well, as shown in Fig. 4, while the latter represents those initially tunneling from the right well to the left well. As shown in Fig. 4, the peak of the occupation probability coherently oscillates between $n_{\rm L}=0$ and $n_{\rm L}=N$ without damping. Due to this coherent oscillation the LGI is violated, as shown in Fig. 4.
The violation of the LGI in Fig. 4 closely resembles the one presented in Fig. 2(a) in Ref. Lambert . This is due to the special coarse graining of the results as the sign ($S_{z}$). Our work represents a concrete realization of the proposal of Lambert et al. for the coarse-grained implementation of LGI violations.
@@
Focusing on the first violation in Fig. 4, this behavior can be explained as follows: The LGI holds until $C_{13}$ starts to decrease, since $C_{12}=C_{13}=C_{23}=1$. The LGI begins to be first violated when $C_{13}$ starts decreasing at $\tau J/2\pi\simeq 0.125$. Since $C_{13}$ abruptly becomes $-1$ while $C_{12}=1$ and $C_{23}$ is still positive, the LGI becomes violated quite abruptly. The value of the peak of $K$ becomes larger with the increase of $N$, due to a more abrupt decrease of $C_{13}$ (see SM).
@@
The persistence of this violation is attributed to the bosons tunneling backwards in Fig. 4 induced by the measurement, which corresponds to the second line of Eq. (Violation of the Leggett-Garg inequality for dynamics of a Bose-Einstein condensate in a double-well potential). In the case of $C_{23}$ the bosons tunneling backwards impede the decrease of $C_{23}$, while such an effect is absent in the case of $C_{13}$, where no measurement is performed at $t_{2}$ [see Fig. 4]. The LGI again begins to hold approximately when $C_{12}$ becomes zero at $\tau J/2\pi\simeq 0.25$ [see Fig. 4].
@@
We also test the LGI for all different initial occupation numbers of the left well $n_{\rm L}(0)$ with $N=100$. In Fig. 4, it can be seen that the peak value of $K$ increases with a larger initial population imbalance (see SM).
Experimental realization. The experimental setup in Ref. Albiez provides a concrete realization of the theoretical predictions in our paper. Phase contrast imaging (PCI) is a promising candidate for implementing our measurement scheme as a nondestructive measurement of a BEC in a double-well potential Andrews ; Bradley ; Higbie ; Ilo-Okeke . PCI can perform measurements in the number basis |$\ket{n_{\rm L}}\ket{n_{\rm R}}$, enabling the measurement of the population in each well nondestructively without atom loss. This approach allows for the acquisition of each correlation function.
@@
Conclusion and discussion. In this Letter, we study violations of the LGI for bosons in a double-well potential. We find that the LGI is violated due to Josephson oscillations in the weak-coupling regime, while it remains unviolated in the strong-coupling regime for $\Lambda$ beyond the critical $\Lambda_{\rm c}$, which depends on the particle number $N$. $\Lambda_{\rm c}$ approaches, but remains below, the critical threshold for the onset of self-trapping, $\Lambda_{\rm st}^{\rm GP}=2$, as derived from semiclassical analysis. It indicates that the nonviolation of the LGI is attributed to the self-trapping phenomena. Remarkably, we observe that the violation of the LGI becomes increasingly significant as the particle number increases. We also find that, in this system, an on-site interaction induces a dephasing effect that destroys temporal correlations, resulting in the LGI no longer being violated afterwards. Furthermore, we have revealed that the LGI is violated for nearly the same parameters for which the Josephson oscillations are observed in Ref. Albiez , thus indicating the presence of macroscopic quantum coherence in the semiclassical regime.
@@
Acknowledgments. We thank anonymous reviewers for insightful comments. We thank G. Kimura, M. Kunimi, and T. Nikuni for fruitful discussions. S.T. thanks F. Dalfovo, G. Ferrari, and A. Recati for useful discussions. R.Y. is sup- ported by the Japanese Society for the Promotion of Science Grant-in-Aid for Scientific Research (KAKENHI, Grants No. 19K14616 and No. 20H01838) and of the WPI program “Sustainability with Knotted Chiral Meta Matter (SKCM2)” at Hiroshima University. S.T. is supported by the Japan Society for the Promotion of Science Grant-in-Aid for Scientific Research (KAKENHI Grant No. 19K03691).